OpenStudy (anonymous):

A batsman deflects a ball by an angle of 45° without changing its initial speed which is equal to 54 km/h. What is the impulse imparted to the ball? (Mass of the ball is 0.15 kg.)

OpenStudy (anonymous):

@Vaidehi09

OpenStudy (anonymous):

umm..is the answer -2.25 N-s ?

OpenStudy (anonymous):

it should be 4.16 kg m/s

OpenStudy (dls):

Impulse = 2 × 0.15 × 15 cos 22.5° = 4.16 kg m/s

OpenStudy (anonymous):

Hw did u get that

OpenStudy (anonymous):

hw that 2 came

OpenStudy (dls):

AO = Incident path OB = Deflected path ∠AOB = Angle between the incident and deflected paths of the ball = 45° ∠AOP = ∠BOP = 22.5° = θ got this?

OpenStudy (anonymous):

isn't impulse = m(v2 - v1) ?

OpenStudy (dls):

Now,according to the question, Initial and final velocities of the ball = v right? horizontal component of the initial velocity = vcos θ along RO vertical component of the initial velocity = vsin θ along PO horizontal component of the final velocity = vcos θ along OS vertical component of the final velocity = vsin θ along OP

OpenStudy (anonymous):

ok...thxxx

OpenStudy (dls):

@Vaidehi09 ∴Impulse imparted to the ball = Change in the linear momentum of the ball so,mv cos theta = (-mvcos theta) =>2mv cos theta

OpenStudy (dls):

the horizontal components of velocities suffer no change the vertical components of velocities are in the opposite directions.

OpenStudy (anonymous):

@DLS but what if we take the initial velocity as x-axis...then wouldn't the diag look like this? what's wrong with this?

OpenStudy (dls):

 if u take x axis

OpenStudy (anonymous):